OpenStudy (anonymous):

Is my answer correct? I'm suppose to use rational exponents to write ^4sqrt x * ^5sqrt 3x is this correct? ^20sqrt 3x^9

OpenStudy (anonymous):

when you type ^4sqrt do you mean:\[\sqrt[4]{x}\]

OpenStudy (anonymous):

\[^4\sqrt x * ^5\sqrt3x\]

OpenStudy (anonymous):

ok so what I did first off was change the sqaure roots into fractions so: x^(1/4) * (3x)^(1/5)

OpenStudy (anonymous):

the 5 is suppose to be an exponent of the ^5sqrt 3x sorry I cant get it to look like the problem

OpenStudy (anonymous):

so it would be (sqrt(3x))^5

OpenStudy (anonymous):

which would change to (3x)^(5/2)

OpenStudy (anonymous):

so than you would have x^(1/4) * (3x)^(5/2) this probably will help you solve it easier if you change the sqaure roots into fractions.

OpenStudy (anonymous):

ok. It still looks confusing to me lol. I feel like I'm learning a foreign language

OpenStudy (anonymous):

yep math can seem that way

OpenStudy (anonymous):

\[\sqrt[4]{x} *(\sqrt{3x})^{5}\] is this what it looks like?

OpenStudy (anonymous):

the first part is right. the second has no () and the ^5 is before the sqrt

OpenStudy (anonymous):

I am changing the 4th root into a fraction by taking the number of times the x is under the radical, and dividing that number by the root

jimthompson5910 (jim_thompson5910):

\[\left(\sqrt[4]{x}\right)\left(\sqrt[5]{3x}\right)\] \[\left(x^{\frac{1}{4}}\right)\left((3x)^{\frac{1}{5}}\right)\] \[\left(x^{\frac{5}{20}}\right)\left((3x)^{\frac{4}{20}}\right)\] \[\left(x^5\cdot(3x)^4\right)^{\frac{1}{20}}\] \[\left(x^5\cdot81x^4\right)^{\frac{1}{20}}\] \[\left(81x^9\right)^{\frac{1}{20}}\] \[\sqrt[20]{81x^9}\]